# Controlling Large Language Model Agents with Entropic Activation Steering

### v1

@@ Abstract
The generality of pretrained large language models (LLMs) has prompted increasing interest in their use as in-context learning agents. To be successful, such agents must form beliefs about how to achieve their goals based on limited interaction with their environment, resulting in uncertainty about the best action to take at each step. In this paper, we study how LLM agents form and act on these beliefs by conducting experiments in controlled sequential decision-making tasks. To begin, we find that LLM agents are overconfident: They draw strong conclusions about what to do based on insufficient evidence, resulting in inadequately explorative behavior. We dig deeper into this phenomenon and show how it emerges from a collapse in the entropy of the action distribution implied by sampling from the LLM. We then demonstrate that existing token-level sampling techniques are by themselves insufficient to make the agent explore more. Motivated by this fact, we introduce Entropic Activation Steering (EAST), an activation steering method for in-context LLM agents. EAST computes a steering vector as an entropy-weighted combination of representations, and uses it to manipulate an LLM agent’s uncertainty over actions by intervening on its activations during the forward pass. We show that EAST can reliably increase the entropy in an LLM agent’s actions, causing more explorative behavior to emerge. Finally, EAST modifies the subjective uncertainty an LLM agent expresses, paving the way to interpreting and controlling how LLM agents represent uncertainty about their decisions.
@@ 1 Introduction
However, in contrast to well-studied decision-making algorithms based on reinforcement learning [25], relatively little is known about how LLM agents come to their decisions through interaction. While the LLM operates at the token level, playing the role of the reasoning engine behind the agent, decisions happen at a higher level of abstraction, after the output text produced by the LLM is parsed into an action. Overall, the interaction between these two levels is not well understood, and it plays a vital role in determining how the agent’s beliefs shape its action distribution.
Indeed, recent work has shown that this process frequently goes awry, causing in-context LLM agents to fail to produce sensible exploratory behavior [10]. They tend to be overconfident, rapidly reducing the uncertainty about their decisions and committing to a particular solution, even when it should be clear that more information is needed. How can we effectively intervene on this behavior?
In this paper, we introduce Entropic Activation Steering (EAST), a method to alter an LLM agent’s subjective uncertainty over decisions and entropy over actions. EAST uses a dataset of logged interactions between the LLM agent and an environment to obtain a steering vector. This vector is computed as an entropy-weighted average of the (run-centered) representations that an LLM produces right before making a decision. Similarly to previous work in activation addition [21], the steering vector is applied at decision time by adding it, at a specific layer, to the representation corresponding to the tokens that are being generated by the LLM.
@@
EAST directly controls the entropy of its distribution over actions, well beyond what is achievable by simply modifying an LLM’s token sampling temperature. Moreover, EAST modifies the subjective uncertainty expressed by an LLM agent in its ReAct-style thoughts [33], towards a less exploitative and more information-seeking attitude. With controlled experiments in bandit tasks expressed in language, we show that EAST is able to steer the agent towards more explorative behavior, effectively addressing the overconfidence exhibited by LLM agents.
We demonstrate that EAST generalizes to variations in prompts and LLMs. Surprisingly, we show that the steering vectors we construct can transfer between tasks which are presented as different natural language scenarios, but equivalent from the sequential decision-making standpoint. Overall, the effectiveness of EAST and our in-depth analyses suggest that LLMs possess an abstract representation of the uncertainty over their decisions, and that it is possible to exercise direct control on it, paving the way to more interpretable and controllable LLM agents.
@@ 2 Background
Modern language models interact with text input through a process of tokenization, in which a body of text is broken down into small units known as tokens [16]. To begin, let $\Omega$ be a finite set of natural language tokens. We consider the set of token sequences of finite length, $\Omega^{*}$, consisting of elements $\omega=(\omega_{1},\ldots,\omega_{n})$ where $n$ is the length of that sequence.
@@
Our experiments focus on a Gaussian multi-armed bandit setting [12], in which the action space $\mathcal{A}$ is a set of possible arm choices and the feedback $F_{t}$ is a string describing a numerical reward drawn from a Gaussian distribution $\mathcal{N}(\mu_{a},\sigma_{a})$ associated to a particular arm $a\in\mathcal{A}$. At each round, the agent has to choose which arm to pick. The task description $P_{0}$ tasks the agent with maximizing the sum of the rewards it receives over time. This setting captures the essential elements of self-evaluation and in-context learning across turns of interaction [23], making them easier to analyze.
@@ 3 Related Work
By studying how LLM agents represent uncertainty and presenting a steering technique specific for agents, our paper connects recent work in LLM agents and in representation engineering [35]. We will now provide an overview of the most relevant work from these two research communities.
LLM-based agents.   LLMs have been recently employed for creating agents, leveraging their capabilities such as proposing actions [31], generating code [29, 15], or evaluating outcomes [9, 11]. In this paper, we focus on in-context LLM agents, which use the ability of LLMs to learn from data in their prompt [1] to process a history of interactions with an environment. We employ an LLM agent multi-armed bandit setup [10, 20]. The advantage of this setup resides in its ability to capture, in a more controlled setting, essential aspects of good decision-making. These systems are typically based on repeated interactions with a task, and heavily rely on the in-context learning abilities of existing LLMs [23, 14, 19]. An important component in our discussion is the relationship between the token generation process and the action extraction process, which is encountered in recent work using reinforcement learning to train LLMs in decision-making tasks [34].
Representations of LLMs and activation steering.   Our analyses of the representation space of LLM agents and our EAST method are closely related to recently proposed techniques for activation steering [24, 26, 21, 13, 32] and, more broadly, to the recent interest in interpreting the activations of LLMs [35, 3]. In particular, similarly to [21], we apply a steering vector during autoregressive unrolling by adding it to the activations at each position of generated tokens. Differently from these methods, the method we will present focuses on a sequential decision-making setting. Furthermore, we intervene on the action entropy of an LLM agent by leveraging a continuous-valued signal instead of the discrete contrastive approach applied in other recent work [21, 26]. Our work is related to recent efforts on the mechanistic interpretability of agents using reinforcement learning to navigate gridworlds [18], or imitating humans to play chess [8]. We instead focus on in-context LLM agents based on pretrained models, connecting recent analyses of the representation space of LLMs in a supervised in-context learning setting [4] to agentic use cases.
@@ 4 A closer look at the uncertainty over actions of an LLM agent
Experimental setting   Following previous work [10], we employ two-armed Gaussian bandits with different means $\mu_{0},\mu_{1}$, which we vary depending on the experiment, and variances $\sigma_{0}=10,\sigma_{1}=10$ unless otherwise specified. We describe the task to the agent with the prompt in Prompt 1, reported in the appendix (which also reports examples of interactions), in which the two arms are described to the agent as Buttons that it can press. The agent is instructed to evaluate both options in order to maximize its score over time. We use the ReAct prompting [33] strategy, which asks the LLM to produce a thought before selecting a particular action. In addition to increasing the reliability of the agent at generating valid actions, inspecting thoughts will also allow us to qualitatively inspect the agent’s expression of its subjective uncertainty. For each round of interaction, we generate $25$ different completions, parse actions from them, and randomly sample from the valid actions. When estimating the entropy of the action distribution, we consider the set of these valid actions. We study LLMs based on the Transformer architecture [28]. We focus on Mixtral-8x7B [5], and report results on DBRX [2] in the appendix. In all cases, the agent-environment interaction is implemented as a dialogue, and we correspondingly use the instruction-tuned versions of these models. Each error bar displayed in the paper shows a bootstrapped 95% confidence interval around the mean, computed using the default behavior of the seaborn python library [30].
@@ 4.1 The behavior of in-context LLM agents in bandit tasks
Previous work [10] has established that a common failure case of current in-context LLM agents comes from overconfident behavior. In the context of a bandit, this overconfidence corresponds to the agent committing to a particular action without sufficient evidence that that particular action is the best one (i.e. leading to a higher expected reward).
@@
To provide another perspective on this phenomenon, Figure 2 (right) shows the evolution of the entropy of the agent’s action distribution as time passes, averaged over the different runs. For all the different configurations, the entropy of the LLM agent’s action distribution rapidly decreases over time, resulting in insufficient exploration of the available options.
@@ 4.2 Connecting token and action generation
In particular, we observe the evolution of the probability of selecting the first button in two steps far in time (step 4 and step 17) in an example run. While in early steps (see step 4) individual tokens in the LLM’s thought progressively determine the action, in later steps (see step 17) the decrease in entropy highlighted in Figure 2 is associated with the evolution of the thought having no effect on the agent’s ultimate decision. Echoing previous work on different forms of chain-of-thought prompting [27], the example shows that a model does not necessarily come to a conclusion at the end of the thought, and that the thought acts as a manifestation of an underlying computational process happening in the representation space, but not always as the only guide to a model’s final decision.
Figure 4: Distribution of choices over two actions (0 and 1) taken by the LLM agent over time when varying the sampling temperature. Increasing temperature until the point at which no action can be parsed from the LLM’s generations does not significantly change the entropy in action distribution.
Having seen the connection between the token-generation and the action-generation processes, it is natural to ask how much intervening on the former can influence the latter, and whether an intervention can counteract the tendency of the LLM agent to be overconfident. The most direct strategy to try to increase the entropy in the generated actions $\pi(\cdot\mid P)$ is to manipulate the entropy in the generated tokens, which is typically achieved by increasing the temperature used during sampling. In Figure 4, we visualize the distribution of agent behaviors on the equal means environment as measured in different runs for various values of sampling temperature, progressively increasing it up to the point in which no valid action can be extracted from any model generation. The results show that temperature does not significantly change the tendency of the model to overcommit, until no run can be completed. This shows that, due to the nature of their interaction, increasing entropy in token generation does not increase the entropy in the action distribution.
@@ 5 Entropic Activation Steering
In the previous section, we have shown that changing the token sampling temperature does not have a significant effect on the action distribution of the agent. We now introduce Entropic Activation Steering (EAST), a method to directly steer the LLM’s action entropy and subjective uncertainty by intervening on its forward pass. EAST consists of two phases: first, computing a steering vector from a dataset of interactions, and second, using the steering vector to modify the behavior of the agent.
In the first phase, given a dataset of prompts $P_{t}^{k}$ obtained by letting the agent interact for $K$ runs of $T$ timesteps each, we compute the activations $z_{t}^{k}=f^{\ell}(P_{t}^{k})$ by giving a prompt $P_{t}^{k}$ as input, forward-passing the LLM, and extracting the representation corresponding to the last token in the prompt. Then, for each prompt, we estimate the entropy $h_{t}^{k}=-\sum_{a\in\mathcal{A}}\pi(a|P_{t}^{k})\log(\pi(a|P_{t}^{k}))$ of the action distribution, by generating $M$ different completions from the LLM, extracting the corresponding action, and computing the entropy on the sampled actions. In practice, we use $M=25$ and only compute the entropy using completions for which the action is successfully parsed. Then, we compute the steering vector as:
@@
with $Z=\sum_{k=1}^{K}\sum_{t=1}^{T}h_{t}^{k}$ a normalizing constant. The steering vector is an entropy-weighted average of the activations in the dataset, in which each activation is centered around the mean of the corresponding run’s activation. As observed in previous work [7], this centering process makes the method more robust to run-specific differences in the representation. Overall, the first phase extracts a representation whose direction is aligned to the direction that leads, on average, to high entropy. Such a simple aggregation mechanism is motivated by the geometry of the representation space [35, 17].
In the second phase, we apply the steering vector to influence the LLM agent’s behavior. While generating a completion, we add the steering vector $\bm{u}$, at each step, to the representation produced by the model at layer $\ell$ at the position of the last token. This yields a steered representation $\widehat{z}_{i}=z_{i}+\beta\bm{u}$, where $\beta$ is a multiplier determining the amount of steering. Note that, when generating subsequent tokens after having applied the intervention on a previous activation, we keep that previous activation in the modified state until the action is executed.

### v2

@@ Abstract
The rise of large language models (LLMs) has prompted increasing interest in their use as in-context learning agents. At the core of agentic behavior is the capacity for exploration, or the ability to actively gather information about the environment. But how do LLM agents explore, and how can we control their exploratory behaviors? To answer these questions, we take a representation-level perspective, and introduce Entropic Activation Steering (EAST), an activation steering method for in-context LLM agents. Firstly, we demonstrate that EAST can effectively manipulate an LLM agent’s exploration by directly affecting the high-level actions parsed from the outputs of the LLM, in contrast to token-level temperature sampling. Secondly, we reveal how applying this control modulates the uncertainty exhibited in the LLM’s thoughts, guiding the agent towards more exploratory actions. Finally, we demonstrate that the steering vectors obtained by EAST generalize across task variants. In total, these results show that LLM agents explicitly encode uncertainty over their actions in their representation space. Our work paves the way for a new understanding of the functioning of LLM agents and to effective control of their decision-making behaviors.
@@ 1 Introduction
However, in contrast to well-studied decision-making algorithms based on reinforcement learning (Sutton & Barto, 2018), relatively little is known about how LLM agents come to their decisions through interaction. While the LLM operates at the token level, playing the role of the reasoning engine behind the agent, decisions happen at a higher level of abstraction, after the output text produced by the LLM is parsed into an action. Overall, the interaction between these two levels is not well understood, and it plays a vital role in determining how the agent’s beliefs shape its action distribution.
Indeed, recent work has shown that this process frequently goes awry, causing in-context LLM agents to fail to produce sensible exploratory behavior (Krishnamurthy et al., 2024). They tend to be overconfident, rapidly reducing the uncertainty about their decisions and committing to a particular solution, even when it should be clear that more information is needed. How can we effectively intervene on this behavior?
In this paper, we introduce Entropic Activation Steering (EAST), a method to alter an LLM agent’s subjective uncertainty over decisions and entropy over actions. EAST uses a dataset of logged interactions between the LLM agent and an environment to obtain a steering vector. This vector is computed as an entropy-weighted average of the (run-centered) representations that an LLM produces right before making a decision. Similarly to previous work in activation addition (Rimsky et al., 2023), the steering vector is applied at decision time by adding it, at a specific layer, to the representation corresponding to the tokens that are being generated by the LLM.
@@
EAST directly controls the entropy of the agent’s distribution over actions, well beyond what is achievable by simply modifying an LLM’s token sampling temperature. Moreover, EAST modifies the subjective uncertainty expressed by an LLM agent in its ReAct-style thoughts (Yao et al., 2022), towards a less exploitative and more information-seeking attitude. With controlled experiments in bandit tasks expressed in language, we show that EAST is able to steer the agent towards more explorative behavior, effectively addressing the overconfidence exhibited by LLM agents.
We demonstrate that EAST generalizes to variations in prompts and LLMs. Surprisingly, we show that the steering vectors we construct can transfer between tasks which are presented as different natural language scenarios, but are equivalent from the sequential decision-making standpoint. Overall, the effectiveness of EAST and our in-depth analyses suggest that LLMs possess an abstract representation of the uncertainty over their decisions, and that it is possible to exercise direct control on it, paving the way to more interpretable and controllable LLM agents.
@@ 2 Background
Modern language models interact with text input through a process of tokenization, in which a body of text is broken down into small units known as tokens (Mielke et al., 2021). To begin, let $\Omega$ be a finite set of natural language tokens. We consider the set of token sequences of finite length, $\Omega^{*}$, consisting of elements $\omega=(\omega_{1},\ldots,\omega_{n})$ where $n$ is the length of that sequence.
@@
Our experiments focus on a Gaussian multi-armed bandit setting (Lattimore & Szepesvari, 2017), in which the action space $\mathcal{A}$ is a set of possible arm choices and the feedback $F_{t}$ is a string describing a numerical reward drawn from a Gaussian distribution $\mathcal{N}(\mu_{a},\sigma_{a})$ associated to a particular arm $a\in\mathcal{A}$. At each round, the agent has to choose which arm to pick. The task description $P_{0}$ tasks the agent with maximizing the sum of the rewards it receives over time. This setting captures the essential elements of self-evaluation and in-context learning across turns of interaction (Shinn et al., 2023), making them easier to analyze.
@@ 3 Related Work
By studying how LLM agents represent uncertainty and presenting a steering technique specific for agents, our paper connects recent work in LLM agents and in representation engineering (Zou et al., 2023). We will now provide an overview of the most relevant work from these two research communities.
LLM-based agents.   LLMs have been recently employed for creating agents, leveraging their capabilities such as proposing actions (Wu et al., 2023), generating code (Wang et al., 2024; Ma et al., 2024), or evaluating outcomes (Klissarov et al., 2024; Kwon et al., 2023). In this paper, we focus on in-context LLM agents, which use the ability of LLMs to learn from data in their prompt (Brown et al., 2020) to process a history of interactions with an environment. We employ an LLM agent multi-armed bandit setup (Krishnamurthy et al., 2024; Park et al., 2024; Schubert et al., 2024; Binz & Schulz, 2023). The advantage of this setup resides in its ability to capture, in a more controlled setting, essential aspects of good decision-making. These systems are typically based on repeated interactions with a task, and heavily rely on the in-context learning abilities of existing LLMs (Shinn et al., 2023; Liu et al., 2023; Mirchandani et al., 2023). An important component in our discussion is the relationship between the token generation process and the action extraction process, which is encountered in recent work using reinforcement learning to train LLMs in decision-making tasks (Zhou et al., 2024).
Representations of LLMs and activation steering.   Our analyses of the representation space of LLM agents and our EAST method are closely related to recently proposed techniques for activation steering (Subramani et al., 2022; Turner et al., 2023; Rimsky et al., 2023; Li et al., 2023; Wu et al., 2024) and, more broadly, to the recent interest in interpreting the activations of LLMs (Zou et al., 2023; Heimersheim & Nanda, 2024). In particular, similarly to (Rimsky et al., 2023), we apply a steering vector during autoregressive unrolling by adding it to the activations at each position of generated tokens. Differently from these methods, the method we will present focuses on a sequential decision-making setting. Furthermore, we intervene on the action entropy of an LLM agent by leveraging a continuous-valued signal instead of the discrete contrastive approach applied in other recent work (Rimsky et al., 2023; Turner et al., 2023). Our work is related to recent efforts on the mechanistic interpretability of agents using reinforcement learning to navigate gridworlds (Mini et al., 2023), or imitating humans to play chess (Karvonen, 2024). We instead focus on in-context LLM agents based on pretrained models, connecting recent analyses of the representation space of LLMs in a supervised in-context learning setting (Hendel et al., 2023) to agentic use cases.
@@ 4 A closer look at the uncertainty over actions of an LLM agent
Experimental setting   Following previous work (Krishnamurthy et al., 2024), we consider two-armed Gaussian bandits with different means $\mu_{0},\mu_{1}$, which we vary depending on the experiment. For ease of analysis, we keep the variances common and fixed to $\sigma_{0}=10,\sigma_{1}=10$ unless otherwise specified. We describe the task to the agent with the prompt in Prompt 1, reported in the appendix (which also reports examples of interactions), in which the two arms are described to the agent as Buttons that it can press. The agent is instructed to evaluate both options in order to maximize its score over time. We use the ReAct prompting (Yao et al., 2022) strategy, which asks the LLM to produce a thought before selecting a particular action. In addition to increasing the reliability of the agent at generating valid actions, inspecting thoughts will also allow us to qualitatively inspect the agent’s expression of its subjective uncertainty. For each round of interaction, we generate $25$ different completions, parse actions from them, and randomly sample from the valid actions. When estimating the entropy of the action distribution, we consider the set of these valid actions. We study LLMs based on the Transformer architecture (Vaswani et al., 2017). We focus on Mixtral-8x7B (Jiang et al., 2024), and also report results for DBRX (Databricks, 2024) in Section 6.3. In all cases, the agent-environment interaction is implemented as a dialogue, and we correspondingly use the instruction-tuned versions of these models. Each error bar displayed in the paper shows a bootstrapped 95% confidence interval around the mean, computed using the default behavior of the seaborn python library (Waskom, 2021).
@@ 4.1 The behavior of in-context LLM agents in bandit tasks
Previous work (Krishnamurthy et al., 2024) has established that a common failure case of current in-context LLM agents comes from overconfident behavior. In the context of a bandit, this overconfidence corresponds to the agent committing to a particular action without sufficient evidence that that particular action is the best one (i.e. leading to a higher expected reward).
@@
To provide another perspective on this phenomenon, Figure 2 (right) shows the evolution of the entropy of the agent’s action distribution $H(\pi(\cdot\mid P_{t}))$ as time passes, averaged over the different runs. For all the different configurations, the entropy of the LLM agent’s action distribution rapidly decreases over time, resulting in insufficient exploration of the available options.
@@ 4.2 Connecting token and action generation
Figure 4: Distribution of choices over two actions (0 and 1) taken by the LLM agent over time when varying the sampling temperature. A darker color corresponds to a more common behavior, and incomplete lines are due to the episode terminating early because of invalid actions. Increasing temperature until the point at which no action can be parsed from the LLM’s generations does not significantly change the entropy in action distribution.
In particular, we observe the evolution of the probability of selecting the first button in two steps far in time (step 4 and step 17) in an example run. While in early steps (see step 4) individual tokens in the LLM’s thought progressively determine the action, in later steps (see step 17) the decrease in entropy highlighted in Figure 2 is associated with the evolution of the thought having no effect on the agent’s ultimate decision. Echoing previous work that has been done on different forms of chain-of-thought prompting (Turpin et al., 2024), the example shows that a model does not necessarily come to a conclusion at the end of the thought, and that the thought acts as a manifestation of an underlying computational process happening in the representation space, but not always as the only guide to a model’s final decision.
Having seen the connection between the token-generation and the action-generation processes, it is natural to ask how much intervening on the former can influence the latter, and whether an intervention can counteract the tendency of the LLM agent to be overconfident. The most direct strategy to try to increase the entropy in the generated actions $\pi(\cdot\mid P)$ is to manipulate the entropy in the generated tokens, which is typically achieved by increasing the temperature used during sampling. In Figure 4, we visualize the distribution of agent behaviors on the equal means environment, across runs, for various values of sampling temperature, progressively increasing it up to the point at which the model fails to consistently produce completions from which a valid action can be parsed. The results show that temperature does not significantly change the tendency of the model to overcommit, until no run can be completed. This shows that, due to the nature of their interaction, increasing entropy in token generation does not increase the entropy in the action distribution.
@@ 5 Entropic Activation Steering
In the previous section, we have shown that changing the token sampling temperature does not have a significant effect on the action distribution of the agent. Now, our motivation is to 1) identify at a mechanistic level what drives exploratory behavior for an LLM agent and 2) develop new controls on this behavior.
A natural candidate for doing this is the class of recent methods based on activation steering, which derive steering vectors from datasets of LLM representations which are iteratively added to the model’s activations during language model generation. These steering vectors are able to modulate complex concepts such as refusal, sycophancy, or hallucination in an LLM’s outputs; the success of this intervention also suggests that the steering vector directions represent semantically meaningful concepts in LLM activation space (Rimsky et al., 2023; Arditi et al., 2024; Turner et al., 2023).
However, existing activation steering techniques are insufficient for intervening on the entropy of the action distribution of an LLM agent, for two main reasons.
  1. 1.
    They typically assume access to a dataset with discrete labels for each prompt e.g. “harmful” or “safe”. In our setting, we are instead interested in controlling a continuous variable.
  2. 2.
    They are designed for non-agentic settings, in which each prompt is an i.i.d. sample from a distribution and there is no feedback loop of interaction with an external system.
To overcome these shortcomings, we now introduce Entropic Activation Steering (EAST), an activation steering method that directly controls the LLM’s action entropy and subjective uncertainty by intervening on its forward pass. EAST consists of two phases: first, computing a steering vector from a dataset of interactions, and second, using the steering vector to modify the behavior of the agent.
In the first phase, given a dataset of prompts $P_{t}^{k}$ obtained by letting the agent interact for $K$ runs of $T$ timesteps each, we compute the activations $z_{t}^{k}=f^{\ell}(P_{t}^{k})$ by giving a prompt $P_{t}^{k}$ as input, forward-passing the LLM, and extracting the $-\ell$ representation corresponding to the last token in the prompt. Then, for each prompt, we estimate the entropy $h_{t}^{k}=-\sum_{a\in\mathcal{A}}\pi(a|P_{t}^{k})\log(\pi(a|P_{t}^{k}))$ of the action distribution, by generating $M$ different completions from the LLM, extracting the corresponding action, and computing the entropy on the sampled actions. In practice, we use $M=25$ and only compute the entropy using completions for which the action is successfully parsed. Then, we compute the steering vector as:
@@
with $Z=\sum_{k=1}^{K}\sum_{t=1}^{T}h_{t}^{k}$ a normalizing constant. The steering vector is an entropy-weighted average of the activations in the dataset, in which each activation is centered around the mean of the corresponding run’s activations.
The use of an entropy weight $h_{t}^{k}$ generalizes the process for steering vector computation proposed in previous work to a continuous setting in the following way. Existing methods typically work in a contrastive fashion, obtaining the steering vector by subtracting the average representation of positive examples from the average representation of negative examples. We can formally see this process as one of defining a weight vector $w$ and then taking an average of the representations in the dataset weighted according to $w$. Through this lens, we can view existing contrastive methods as assigning components $w_{i}$ such that $w_{i}\in\{-1,1\}$ according to the label in the dataset. In this perspective, one can interpret the entropy weight $h_{t}^{k}$ as a continuous target, representing an extension of the implicit weighting implied by existing methods.
To handle the interactive nature of the decision task, EAST aggregates over independent runs $P^{k}$ and timesteps within those runs $P_{t}^{k}$. In contrast to the naïve approach of simply summing the corresponding activations, EAST normalizes each activation $z_{t}^{k}$ within a run by the average activation over that run. We found in preliminary experiments that this approach is essential to produce functioning steering vectors. This suggests that as the interaction history encoded in the prompt grows, LLM representations specialize to the particular events of that history; our normalization method works to target the specific component of representation space responsible for explorative behavior that is common across runs.
Overall, the first phase extracts a representation whose direction is aligned with the direction that leads, on average, to high entropy. In the second phase, we apply the steering vector to influence the LLM agent’s behavior. While generating a completion, we add the steering vector $\bm{u}$, at each step, to the representation produced by the model at layer $\ell$ at the position of the last token. This yields a steered representation $\widehat{z}_{i}=z_{i}+\beta\bm{u}$, where $\beta$ is a multiplier determining the amount of steering. Note that, when generating subsequent tokens after having applied the intervention on a previous activation, we keep that previous activation in the modified state until the action is executed.